OpenStudy (anonymous):

Solve the following system of equations. Variables x,y,z. 7x-y+z=-22 3x+2y-3z=-27 x-3y+2z=1

OpenStudy (jmark):

Given 7x-y+z=-22.........(1) 3x+2y-3z=-27 …........(2) x-3y+2z=1 …........(3) multiply by 3 in equation (1) and add with equation (2) =>21x-3y+3z=-66........(1) =>3x+2y-3z=-27 ..--------------------- ….24x-y=-93.......(4) multiply by 3 in equation (3) and multiply by 2 in equation (2) then add =>6x+4y-6z=-54 =>3x-9y+6z=3 ...------------------- ….9x-5y=-51.......(5) multiply by 5 in equation (4) then subtract with equation (5) =>120x-5y=-465 =>9x-5y=-51 ….-....+....+ ..---------------- 111x+0= -414 =>x= -138/37 =>y=129/37 =>z=281/37 See also Parametric Equation of a Circle http://goo.gl/Va1n67

OpenStudy (anonymous):